OpenStudy (anonymous):

*WILL MEDAL AND FAN* The speed of light is 3 × 10^8 meters per second. What is this speed in standard form? A.3,000,000,000 meters per second B.300,000,000 meters per second C.30,000,000 meters per second D.300,000 meters per second

OpenStudy (anonymous):

3 x 10^8

OpenStudy (anonymous):

B.300,000,000 meters per second

OpenStudy (anonymous):

that's the point, baby :)

OpenStudy (anonymous):

Thank you! can you help me with couple more

OpenStudy (anonymous):

of course, go ahead

OpenStudy (anonymous):

day is equivalent to 8.64 × 10^4 seconds. What is this time in standard form?

OpenStudy (anonymous):

8.64 x 10^4

OpenStudy (anonymous):

86,400 seconds

OpenStudy (anonymous):

Ahh Thank you :)

OpenStudy (anonymous):

you are welcome.., :)

OpenStudy (anonymous):

im sorry if im bothering you. if you dont want me to show you the rest let me know

OpenStudy (anonymous):

No problem. I'd love to help you, trust me

OpenStudy (anonymous):

You are not bothering me .., ;)

OpenStudy (anonymous):

Which of the following best represents the speed of Halley's comet when it is nearest to Earth?

OpenStudy (anonymous):

ok, hold on

OpenStudy (anonymous):

1 × 10-5 kilometers per year 1 × 105 kilometers per year 1 × 105 kilometers per hour 1 × 10-5 kilometers per hour

OpenStudy (anonymous):

all of them are exponets and also all of them are times

OpenStudy (anonymous):

I think the answer is 1 × 10^5 kilometers per hour

OpenStudy (anonymous):

1 x 10^5 kilometers per hour

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

The length of an object is measured as 45.7 centimeters. What is this measurement in scientific notation?

OpenStudy (anonymous):

yeay! :)

OpenStudy (anonymous):

4.57 × 101 centimeters 0.457 × 102 centimeters 457 × 101 centimeters 4,570 × 102 centimeters

OpenStudy (anonymous):

Also same there all exponets and all of them are times

OpenStudy (anonymous):

4.57 x 10^1 centimeters

OpenStudy (anonymous):

of course!

OpenStudy (anonymous):

3 more :)

OpenStudy (anonymous):

the answer is A

OpenStudy (anonymous):

The equatorial radius of Earth is approximately 6 × 103 km, while the equatorial radius of Saturn is approximately 6 × 104 km. Which of the following is true? The equatorial radius of Earth is approximately one hundred times that of Saturn. The equatorial radius of Saturn is approximately one hundred times that of Earth. The equatorial radius of Saturn is approximately ten times that of Earth. The equatorial radius of Earth is approximately ten times that of Saturn.

OpenStudy (anonymous):

ok , please

OpenStudy (anonymous):

6 x 10^3 km and 6 x 10^4 km

OpenStudy (anonymous):

The equatorial radius of Saturn is approximately ten times that of Earth.

OpenStudy (anonymous):

C right?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

what else ??

OpenStudy (anonymous):

The mass of an electron is approximately 9 × 10-28 grams, while the mass of a neutron is approximately 2 × 10-24 grams. Which of the following is true?

OpenStudy (anonymous):

9 x 10^-28 and 2 x 10^-24 grams

OpenStudy (anonymous):

The mass of a neutron is approximately 2,000 times the mass of an electron. The mass of a neutron is approximately 20,000 times the mass of an electron. The mass of a neutron is approximately 10,000 times the mass of an electron. The mass of a neutron is approximately 1,000 times the mass of an electron.

OpenStudy (anonymous):

the mass of a electron = Me = 9 x 10^-28 grams the mass of a neutron = Mn = 2 x 10^-24 grams So \[\frac{ Mn }{ Me } = \frac{ 2 \times 10^{-24} }{ 9 \times 10^{-28}} = \frac{ 2 \times 10^{-24+28} }{ 9 } = \frac{ 2 \times 10^{4} }{ 9 } \] \[Mn = \frac{ 2 }{ 9 } \times 10^{4} \]

OpenStudy (anonymous):

i mean \[Mn = \frac{ 2 }{ 9 } \times 10^{4} Me\]

OpenStudy (anonymous):

So, The mass of a neutron is approximately 10,000 times the mass of an electron.

OpenStudy (anonymous):

Ok got it but just one question whats the answer? im kinda confused

OpenStudy (anonymous):

the answer is The mass of a neutron is approximately 10,000 times the mass of an electron.

OpenStudy (anonymous):

the reason why the mass of a neutron is approximately 10,000 times the mass of an electron because the mass of a neutron is (2/9) x 10^4 mass of an electron, like what i've calculated above

OpenStudy (anonymous):

are you still confused for the last??

OpenStudy (anonymous):

@Kyliejones :)